# Toponium physics at the Large Hadron Collider

### v1

@@ 3 Toponium phenomenology at the LHC
The left panel of figure 1 presents the doubly-differential cross section in $m_{t\bar{t}}$ and $p^{*}$. The distribution’s shape follows directly from the properties of the QCD Green’s function, exhibiting a higher density of events around a binding energy of $-2$ GeV and low $p^{*}$. This behaviour hence reflecting the dependence of the magnitude squared of the Green’s function $|\widetilde{G}(E;p^{*})|^{2}$ on $E$ and $p^{*}$, our results suggest that precise measurements of this observable at the LHC could give direct access to non-perturbative QCD dynamics. The internal structure of toponium, or in other words the heaviest and most compact hadron in the Standard Model, has thus the potential to become experimentally observable.
@@
This conclusion is further supported by the results presented in the left panel of figure 2, which shows the invariant mass distribution of the reconstructed top-antitop system obtained by integrating the two-dimensional distribution of figure 1 over $p^{*}$. We compare predictions from non-relativistic QCD (orange) with those from the two toy model configurations: one including the re-weighting of eq. (4) and a width $\Gamma_{\eta_{t}}=7$ GeV (blue), and another ignoring any re-weighting and using $\Gamma_{\eta_{t}}=2.8$ GeV (green). While the position of the peak is similar in all three cases, which is consistent with the two-dimensional results, the shape of the spectrum differs significantly. Both toy models underestimate the event rate at larger invariant masses, where contributions from higher toponium excitations are expected. This discrepancy must consequently be more pronounced in the narrower width scenario, as found in our predictions.
@@ 4 Conclusion
Including toponium effects in top-antitop theoretical predictions is therefore essential for improving the precision of the Standard Model predictions, especially in light of recent high-statistics LHC measurements that reveal anomalies potentially linked to the omission of these contributions in simulations. Looking forward, future work should focus on quantifying the associated theoretical uncertainties, extending the formalism to account for higher-spin and higher-partial-wave toponium states and matching predictions to higher-order perturbative calculations.
@@ Acknowledgements
This work was supported by the French Agence Nationale de la Recherche (Grant ANR-21-CE31-0013), a Natural Science Basic Research Program of Shaanxi Province in China (2023-JC-YB-041) and the JSPS (KAKENHI Grants No. 21H01077 and 23K03403).

### v2

@@ 3 Toponium phenomenology at the LHC
The left panel of figure 1 presents the doubly-differential cross section in $m_{t\bar{t}}$ and $p^{*}$. The distribution’s shape follows directly from the properties of the QCD Green’s function, exhibiting a higher density of events around a binding energy of $-2$ GeV and low $p^{*}$. This behaviour hence reflects the dependence of the magnitude squared of the Green’s function $|\widetilde{G}(E;p^{*})|^{2}$ on $E$ and $p^{*}$. Our results thus suggest that precise measurements of this observable at the LHC could give direct access to non-perturbative QCD dynamics. The internal structure of toponium, or in other words the heaviest and most compact hadron in the Standard Model, has thus the potential to become experimentally observable.
@@
This is further supported by the results presented in the left panel of figure 2, which shows the invariant mass distribution of the reconstructed top-antitop system obtained by integrating the two-dimensional distribution of figure 1 over $p^{*}$. We compare predictions from non-relativistic QCD (orange) with those from the two toy model configurations: one including the re-weighting of eq. (4) and $\Gamma_{\eta_{t}}=7$ GeV (blue), and another ignoring any re-weighting and using $\Gamma_{\eta_{t}}=2.8$ GeV (green). We observe that all three predictions exhibit a peak consistent with the two-dimensional results and the expectation from the toponium ground states. However, the spectrum shapes differ significantly. In particular, both toy models underestimate the event rate at larger invariant masses, where contributions from higher-spin toponium excitations and octet states should be in order, as shown in [23, 24]. The discrepancy with the toy models is thus more pronounced in the narrower width scenario, and this should push for the inclusion of both higher-spin and colour-octet contributions in the Green’s function.
@@ 4 Conclusion
Including toponium effects in $t\bar{t}$ theoretical predictions is therefore essential for improving the precision of the Standard Model predictions, especially in light of recent high-statistics LHC measurements that reveal anomalies potentially linked to the omission of these contributions in simulations. Looking forward, future work should focus on quantifying the associated theoretical uncertainties, extending the formalism to account for higher-spin and higher-partial-wave toponium states and matching predictions to higher-order perturbative calculations.
@@ Acknowledgements
This work was supported by the Agence Nationale de la Recherche Grant ANR-21-CE31-0013.